A footwear factory sold 600 pairs of slippers. The number sold at market stalls is twice the number sold at department stores,and the number sold by foot salesman is one-sixth the number sold at market stalls. How many pairs of slippers were sold at market stalls?
Answer by mananth(16946)

Let departmental stores sell x numbers
market stalls sell = 2x
foot salesman sells 2x/6
...
Total sold = 600
x+2x+x/3 =600
10x/3 =600
10x=1800
x= 180 sold at departmental stores
market stalls = 2x = 360
foot salesman = x/3 =60
..
180+360+60= 600
...
m.ananth@hotmail.ca